OpenStudy (anonymous):

A box has 1 red marble, 3 blue marbles, and 4 green marbles. Maya draws a blue marble randomly from the box, replaces it, and then draws another marble randomly. What is the probability of drawing 2 blue marbles in a row? Explain your answer.

OpenStudy (anonymous):

@Michele_Laino u think u culd help me some more ^.^?

OpenStudy (michele_laino):

at the first draw, the probability for getting a blue marble, is: 3/8 then Maya replaces it into the box

OpenStudy (michele_laino):

at the second draw the probability to getting a blue marble, is the same, namely 3/8

OpenStudy (michele_laino):

now, what is the requested probability?

OpenStudy (anonymous):

so both probabilities is 3/8 ^.^ ?

OpenStudy (michele_laino):

yes! Since Maya replaces the extracted marble into the box every time

OpenStudy (michele_laino):

then the requested probability is: 3/8 * 3/8 = ...?

OpenStudy (michele_laino):

\[\frac{3}{8} \times \frac{3}{8} = ...?\]

OpenStudy (anonymous):

its 9/64? i think

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

how do i simplify tht ^.^

OpenStudy (michele_laino):

we can not simplify that fraction, since 9 and 64 have no common divisors

OpenStudy (anonymous):

oh okay thx again is it ok if i tag u in a few more questions?

OpenStudy (michele_laino):

ok!